Circular No. 8151

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 L2 (LINEAR) An apparently asteroidal object found by LINEAR (discovery observa-

tion below), and posted on the NEO Confirmation Page, has been found to be cometary on CCD images taken by S. Sanchez, R. Stoss, and J. Nomen (Mallorca, 0.30-m f/9 reflector; 10′′ coma on June 12.95 UT) and by Š. Gajdoš (Modra, 0.6-m f/5.5 reflector; diffuse with coma diameter ∼ 5′′ on June 13.97; m_{1} = 18.0).

2003 UT  α_{2000}  δ_{2000}  m_{2}
June 12.32879  19^{h}30^{m}27.^{s}04  +56^{o}54′17′′.6  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-L55.

T = 2004 Feb. 14.914 TT  ω = 139.^{o}542
Ω = 272.220 }2000.0
q = 2.25126 AU  i = 81.546

SUPERNOVA 2002ic M. Hamuy and M. Phillips, Carnegie Observatories; N. Suntzeff, Cerro

Tololo Inter-American Observatory; and J. Maza, University of Chile, report that spectroscopic observations (range 380–930 nm) of SN 2002ic (cf. IAUC 8019), obtained between 2002 Nov. 29 and 2003 Jan. 27 with the Baade telescope (+ Boller & Chivens spectrograph) and the Dupont telescope (+ WFCCD), show a persistent, strong Hα emission and a weaker Hβ emission. The Hα profile exhibits a double component, one unresolved (FWHM < 300 km/s) and the other with FWHM ∼ 1800 km/s — both of which remained strong over the first two months of evolution. This feature is reminiscent of type-IIn supernovae and suggests that this is the first detection of interaction between the ejecta of a type-Ia supernova (cf. IAUC 8028) and a hydrogen-rich circumstellar medium. Hamuy et al. add that SN 2002ic is about to reappear from behind the sun, so further observations are urged in order to check or rule out this possibility.

SUPERNOVA 2003ga Further to IAUC 8148, R. Ellis et al. report that a spectrum of SN

2003ga (cf. IAUC 8149), obtained on June 2 with the Keck I telescope, shows it to be an early type-II supernova with a featureless blue spectrum (and galaxy redshift 0.29).

2003 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green